OpenStudy (anonymous):

how do you find the angle in a triangle when you are given the area and both sides on the triangle

OpenStudy (anonymous):

A = (1/2) b*h now, it's important to be able to tell if the sides you know actually qualify as base and height. Is one of those sides perpendicular to the other?

OpenStudy (anonymous):

There is an equation\[A = {1 \over 2} ab \sin (\theta)\]where theta is the angle between the sides a and b. Then we have that\[\theta = \sin^{-1} ({2A \over ab})\]

OpenStudy (anonymous):

thakyou for that I really apriciate it

OpenStudy (anonymous):

Yeah sure! By the way, the formula is essentially the same as "one half base times height"; it's just that if you draw a right triangle inside of your triangle you see that a*sin(theta) (or b*sin(theta), depending on how it's labeled) is the height, and the remaining side is the base.

OpenStudy (unklerhaukus):

A triangle is just half a rectangle